Derek will deposit $5,341.00 per year for 9.00 years into an account that earns 8.00%. Assuming the first deposit is made 4.00 years from today, how much will be in the account 34.00 years from today?
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

=5341/1.08^4+5341/1.08^5+...........+5341/1.08^12

=5341[1/1.08^4+1/1.08^5+..............+1/1.08^12]

=(5341*4.95898103)

=26485.92(Approx).

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A =26485.92*(1.08)^34

=$362,595.75(Approx).